# Rigidity of proper quasi-homogeneous domains in positive flag manifolds

### v1

@@ 7.1.2. $\mathscr{R}$-extremal points
Following the notation of [LZ19], we denote by $\operatorname{Extr}_{\mathscr{R}}(\Omega)$ the set of $\mathscr{R}$-extremal points of $\Omega$. We will see in Section 7.2 that this set is never empty. In this section we aim at showing the following link between visuality and $\mathscr{R}$-extremality:
Proposition 7.4.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper quasi-homogeneous domain. Then any $\mathscr{R}$-extremal point of $\Omega$ is visual.
@@
  1. (1)
    Theorem 1.4 will actually imply that the reciprocal of Proposition 7.4 is true.
  2. (2)
    Proposition 7.4 is an analogue of classical Vey’s Lemma (Fact 1.8) of real projective geometry, in the context of Shilov boundaries of Hermitian symmetric spaces. Note that in Proposition 7.4, we assume that $\Omega$ is quasi-homogeneous, which is not needed in Vey’s lemma. The reason lies in the definition of the Kobayashi metric: given two sequences $(x_{n}),(y_{n})$ of $\Omega$ such that $x_{n}\rightarrow p\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$, the minimal integer $N\in\mathbb{N}$ such that $\mathscr{C}_{x_{n},y_{n}}^{N}(\Omega)$ is nonempty could go to infinity while the distance $K_{\Omega}(x_{n},y_{n})$ remains bounded. Then it could be possible for $(y_{n})$ to converge to a point $q\in\partial\Omega\smallsetminus\{p\}$. This phenomenon is avoided when $\Omega$ is quasi-homogeneous.
For the proof of Proposition 7.4, we follow the strategy of [LZ19, Thm 7.4]. We will need the following definition:
@@
The first step in the proof of Proposition 7.4 is a weaker version, where we replace $K_{\Omega}$ by $K^{N}_{\Omega}$ (note that on the other hand we do not need the quasi-homogeneity assumption in the next lemma):
@@
To prove Proposition 7.4, we need to strengthen Lemma 7.7 and replace $K_{\Omega}^{N}$ by $K_{\Omega}$ in the assumptions. Lemma 7.9 below, together with the quasi-homogeneity assumtion, will allow us to do so. To prove Lemma 7.9, we prove the auxiliary Lemma 7.8.
Since $K_{\Omega}$ is continuous, one can define the diameter of a compact subset $\mathscr{C}$ of $\Omega$ as:
$\operatorname{diam}(\mathscr{C}):=\max\left\{K_{\Omega}(x,y)\mid x,y\in\mathscr{C}\right\}.$
Then one has:
@@
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain, and let $D$ be a diamond such that $\overline{D}\subset\Omega$. Then for any $x,y\in D$, there exists an $\mathsf{n}(G)$-chain from $x$ to $y$ with $K_{\Omega}$-length less than or equal to $\mathsf{n}(G)\operatorname{diam}(\overline{D})$. In particular,
$K_{\Omega}^{\mathsf{n}(G)}(x,y)\leq\mathsf{n}(G)\operatorname{diam}(\overline{D}).$
@@
    We know by Lemma 6.2 that there exists an $\mathsf{n}(G)$-chain $\gamma=(x_{0},\cdots,x_{N})$ from $x$ to $y$ contained in $D$. By Equation Proposition 6.8, one has
    $K_{\Omega}^{\mathsf{n}(G)}(x,y)\leq\mathsf{len}_{\Omega}(\gamma)=\sum_{i=0}^{N-1}K_{\Omega}(x_{i},x_{i+1})\leq\sum_{i=0}^{N-1}K_{D}(x_{i},x_{i+1})\leq\mathsf{n}(G)\operatorname{diam}(\overline{D}).$
@@
Lemma 7.9.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain, and let $\mathscr{K}\subset\Omega$ be a compact subset. Then there exist $N\in\mathbb{N}$ and $M>0$ such that for any $x,y\in\mathscr{K}$,
$K_{\Omega}^{N}(x,y)\leq M.$
  • Proof.
    First choose, for any $x\in\mathscr{K}$, a diamond $D_{x}$ of diameter less than 1 for $K_{\Omega}$, and containing $x$. Such a diamond exists by Remark 4.9.(1) and Proposition 6.7.
    For any $x,y\in\mathscr{K}$, there is a minimal $N(x,y)\in\mathbb{N}$ such that
    $M(x,y):=K_{\Omega}^{N(x,y)}(x,y)<+\infty.$
    Let $(z,z^{\prime})\in D_{x}\times D_{y}$. There exists $\gamma\in\mathscr{C}_{x,y}^{N(x,y)}(\Omega)$ such that $\mathsf{len}_{\Omega}(\gamma)\leq M(x,y)+1$. Besides, by Lemma 7.8, there exist $\gamma_{1}\in\mathscr{C}_{z,x}^{\mathsf{n}(G)}(\Omega)$ and $\gamma_{2}\in\mathscr{C}_{y,z}^{\mathsf{n}(G)}(\Omega)$, both of $K_{\Omega}$-length less than or equal to $\mathsf{n}(G)$. If $N_{0}:=N(x,y)+2\mathsf{n}(G)$ and $\gamma_{3}$ is the concatenation of $\gamma_{1}$, $\gamma$ and $\gamma_{2}$, then $\gamma_{3}\in\mathscr{C}_{z,z^{\prime}}^{N_{0}}(\Omega)$ and for any $N\geq N$ one has
    $K^{N}_{\Omega}(z,z^{\prime})\leq K^{N_{0}}_{\Omega}(z,z^{\prime})\leq\mathsf{len}_{\Omega}(\gamma_{3})=\mathsf{len}_{\Omega}(\gamma_{1})+\mathsf{len}_{\Omega}(\gamma)+\mathsf{len}_{\Omega}(\gamma_{2})\leq 2\mathsf{n}(G)+M(x,y)+1.$
    Let $F\subset\mathscr{K}$ be a finite subset such that $\mathscr{K}\subset\bigcup_{x\in F}D_{x}$. Then for all $z,z^{\prime}\in\mathscr{K}$, one has $K_{\Omega}^{N}(z,z^{\prime})\leq M$, where
    $N$ $:=\max\{N(x,y)\mid x,y\in F\}+2\mathsf{n}(G);$
    $M$ $:=\max\{M(x,y)\mid x,y\in F\}+2\mathsf{n}(G)+1.\quad\hfill\square$
  • Proof Proposition 7.4.
    Note that $\Omega$ is proper and dually convex by Proposition 4.4, hence Lemmas 7.7 and 7.9 apply to $\Omega$. Write $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$ where $\mathscr{K}\subset\Omega$ is a compact subset. Let $(x_{k})\in\Omega^{\mathbb{N}}$ be such that $x_{k}\rightarrow p$, and let $(y_{k})\in\Omega^{\mathbb{N}}$ be such that there exists $M>0$ with $K_{\Omega}(x_{k},y_{k})\leq M$ for all $k\in\mathbb{N}$. For all $k\in\mathbb{N}$, there exists $g_{k}\in\mathsf{Aut}(\Omega)$ such that $g_{k}\cdot x_{k}\in\mathscr{K}$. Let
    $\mathscr{K}^{\prime}:=\{x\in\Omega\mid K_{\Omega}(x,\mathscr{K})\leq M\}.$
    The set $\mathscr{K}^{\prime}$ is compact by properness of $K_{\Omega}$ (see Corollary 6.13). Then, by Lemma 7.9, there exist $N\in\mathbb{N}$ and $M^{\prime}>0$ such that $K^{N}_{\Omega}(a,b)\leq M^{\prime}$ for all $a,b\in\mathscr{K}^{\prime}$. By $\mathsf{Aut}(\Omega)$-invariance of $K_{\Omega}$, for any $k\in\mathbb{N}$ one has $g_{k}\cdot x_{k},\ g_{k}\cdot y_{k}\in\mathscr{K}^{\prime}$. Hence by $\mathsf{Aut}(\Omega)$-invariance of $K_{\Omega}^{N}$ one has
    $K_{\Omega}^{N}(x_{k},y_{k})=K_{\Omega}^{N}(g_{k}\cdot x_{k},\ g_{k}\cdot y_{k})\leq M^{\prime}.$
@@
7.2. A special set of $\mathscr{R}$-extremal points: the strongly $\mathscr{R}$-extremal points
@@
    Contrary to the notion of $\mathscr{R}$-extremality, that of strong $\mathscr{R}$-extremality is only defined for a domain $\Omega$ which is proper in $\mathbb{A}$. It is not clear at first that this second notion is invariant under $\mathsf{Aut}(\Omega)$. Using Lemma 3.5 it is actually possible to show that this is the case whenever $\Omega$ is dually convex, although we will not use this fact. We will only prove this invariance in the quasi-homogeneous case; see Lemma 8.2.
@@
In this section we finish the proof of Theorem 1.4. We take $\Omega\subset\mathbf{Sb}(G)$ a quasi-homogeneous domain, proper in $\mathbb{A}$.
Let $x\in\Omega$ and let $p_{0}\in\partial\Omega\cap\mathbf{J}^{-}(x)$ and $q_{0}\in\partial\Omega\cap\mathbf{J}^{+}(x)$ be two strongly $\mathscr{R}$-extremal points of $\partial\Omega$ given by Lemma 7.10. Then in particular $p_{0},q_{0}\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$ (Remark 7.11.(2)), so by Lemma 7.1 and Proposition 7.4, one has
@@
(8.3) $\mathsf{Aut}(\Omega)\leq\mathsf{Aut}(\mathbf{D}(p_{0},q_{0})).$
The next proposition ends the proof of Theorem 1.4.
Proposition 8.3.
We have $\Omega=\mathbf{D}(p_{0},q_{0})$.
  • Proof.
    Since $\Omega$ is open in the domain $D:=\mathbf{D}(p_{0},q_{0})$, it suffices to prove that $\Omega$ is closed in $D$. Let $(x_{k})\in\Omega^{\mathbb{N}}$ converging to a point $x\in D$. Write $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$, with $\mathscr{K}\subset\Omega$ a compact subset. For any $k\in\mathbb{N}$ there exist $(g_{k},z_{k})\in\mathsf{Aut}(\Omega)\times\mathscr{K}$ such that $g_{k}\cdot z_{k}\rightarrow x$. Up to extracting there exists $z\in\mathscr{K}$ such that $z_{k}\rightarrow z$. Then
    $K_{D}(g_{k}\cdot z,x)\leq K_{D}(g_{k}\cdot z,g_{k}\cdot z_{k})+K_{D}(g_{k}\cdot z_{k},x)=K_{D}(z,z_{k})+K_{D}(g_{k}\cdot z_{k},x)\underset{k\rightarrow+\infty}{\longrightarrow}0.$
    The last equality holds by Proposition 6.5.(2) and (8.3), and the limit by Proposition 6.7. Since $K_{D}$ generates the standard topology on $D$ (again by Proposition 6.7), we have $g_{k}\cdot z\rightarrow x$.
[…]
Remark 8.4.
The arguments of this section generalize those of [CG24]. However, the causal convexity arguments used in the latter, and specifically [CG24, Prop. 6.2], do not hold anymore. Instead, in the present paper, causality in $\mathbf{Sb}(G)$ appears in the proofs of Propositions 8.2 and 8.3.
@@ 9. Application: closed manifolds with proper development
    Let $\pi_{1}(M)$ be the fundamental group of $M$ and let $\mathsf{hol}:\pi_{1}(M)\rightarrow G$ be the holonomy of $M$, so that the map $\mathsf{dev}$ is $\mathsf{hol}$-equivariant. Let $\Omega=\mathsf{dev}(\widetilde{M})$. Since $M$ is closed, there is a compact fundamental domain $\mathscr{K}\subset\widetilde{M}$ intersecting every $\pi_{1}(M)$-orbit. By $\mathsf{hol}$-equivariance of $\mathsf{dev}$, the compact set $\mathsf{dev}(\mathscr{K})\subset\Omega$ intersects any $\mathsf{hol}(\pi_{1}(M))$-orbit. Hence $\Omega$ is quasi-homogeneous. Theorem 1.4 implies that $\Omega$ is a diamond and is $\left(H_{1},H_{2}\right)$-equivariantly diffeomorphic to the symmetric space ×$\mathbb{R}\times X_{L_{s}}$, where $H_{1},H_{2}$ are finite-index subgroupds of $\mathsf{Aut}(\Omega),\mathrm{Isom}\left(\mathbb{R}\times X_{L_{s}}\right)$ respectively. Up to taking a finite cover of $M$, we may assume that $H_{1}=\mathsf{Aut}(\Omega)$ and $H_{2}=\mathrm{Isom}(\mathbb{R}\times X_{L_{s}})$.

### v2

@@ 7.1. $\mathscr{R}$-extremal points
Following the notation of [LZ19], we denote by $\operatorname{Extr}_{\mathscr{R}}(\Omega)$ the set of $\mathscr{R}$-extremal points of $\Omega$. We will see in Section 7.2 that this set is never empty.
Whenever $\Omega$ is almost-homogeneous, $\mathscr{R}$-extremal points satisfy a strong geometric property:
Theorem 7.2.
Assume that $\Omega$ is a proper almost-homogeneous domain of $\mathbf{Sb}(G)$. Let $p\in\partial\Omega$ be an $\mathscr{R}$-extremal point. Then $\operatorname{Z}_{p}\cap\Omega=\emptyset$.
@@
Theorem 7.2 is specific to our context: given a point $p$ of an arbitrary flag manifold $\mathscr{F}_{\Theta}$, the set $\operatorname{Z}_{p}$ is a subset of $\mathscr{F}_{\Theta}^{-}$, which cannot be $G$-equivariantly identified with $\mathscr{F}_{\Theta}$ in general. There are versions of Theorem 7.2 for flag manifolds that are not self-opposite, but they express a weaker geometric property for $\mathscr{R}$-extremal points (see e.g. [Gal25]). For instance, for the flag manifold $\mathbb{P}(\mathbb{R}^{n})$ with $n\geq 2$, which is as far as possible from being self-opposite, this weaker version of Theorem 7.2 expresses that any $\mathscr{R}$-extremal point of $\partial\Omega$ is contained in a supporting projective hyperplane to $\Omega$, which is already a consequence of the convexity of $\Omega$.
For the proof of Theorem 7.2, we follow the strategy of [LZ19, Thm 7.4]. We will need the following definition:
@@
We can now prove an analogue of Fact 1.8:
@@
Let $\Omega$ be a proper almost-homogeneous domain. Then for all $p\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$ there exists $(g_{n})\in\mathsf{Aut}(\Omega)^{\mathbb{N}}$ such that for every compact subset $\mathscr{K}\subset\Omega$, one has $g_{n}\cdot\mathscr{K}\rightarrow\{p\}$ for the Hausdorff topology.
@@
    Since $\Omega$ is almost-homogeneous we can find $x\in\Omega$ and some sequence $(g_{k})\in\mathsf{Aut}(\Omega)^{\mathbb{N}}$ such that $g_{k}\cdot x\underset{k\rightarrow+\infty}{\longrightarrow}p.$ Now let $y\in\Omega$ and $N\in\mathbb{N}$ such that $K_{\Omega}^{N}(x,y)<+\infty$. Then, by $\mathsf{Aut}(\Omega)$-invariance of $K_{\Omega}^{N}$, one has
    $K_{\Omega}^{N}(g_{k}\cdot x,g_{k}\cdot y)=K_{\Omega}^{N}(x,y)\quad\forall k\in\mathbb{N}.$
    Thus by Lemma 7.5, we have $g_{k}\cdot y\rightarrow p$. This holds for all $y\in\Omega$.
Let $\mathscr{K}\subset\Omega$ be a compact subset. If the sequence $g_{k}\cdot\mathscr{K}$ does not converge to $\{p\}$ for the Hausdorff topology, then there is a neighborhood $\mathscr{V}$ of $p$ in $\mathbf{Sb}(G)$ and a sequence $(y_{k})\in\mathscr{K}^{\mathbb{N}}$ such that $g_{k}\cdot y_{k}\notin\mathscr{V}$ for all $k\in\mathbb{N}$. Since $\mathscr{K}$ is a compact subset of $\Omega$ by Corollary 6.15, up to extracting we may assume that there exists $y\in\mathscr{K}$ such that $y_{k}\rightarrow y$. Then $(g_{k}\cdot y)$ converges to $p$. But $K_{\Omega}(g_{k}\cdot y_{k},g_{k}\cdot y)\rightarrow 0$, so $g_{k}\cdot y_{k}\rightarrow p$. But this is impossible, since we assumed that $y_{k}\notin\mathscr{V}$ for all $k$.
@@
We can now prove Theorem 7.2:
  • Proof of Theorem 7.2.
    Applying Lemma 7.6 to a compact subset $\mathscr{K}\subset\Omega$ with nonempty interior, and using the KAK decomposition, up to extracting, we may assume that there exists $q\in\mathbf{Sb}(G)$ such that $(g_{k})$ is $(p,q)$-contracting, in the sense that
    $g_{k|\mathbf{Sb}(G)\smallsetminus\operatorname{Z}_{q}}\rightarrow p,$
@@
7.2. Strongly $\mathscr{R}$-extremal points
@@
    Contrary to the notion of $\mathscr{R}$-extremality, that of strong $\mathscr{R}$-extremality is only defined for a domain $\Omega$ which is proper in $\mathbb{A}$. It is not clear at first that this second notion is invariant under $\mathsf{Aut}(\Omega)$. Using Lemma 3.5 it is actually possible to show that this is the case whenever $\Omega$ is dually convex, although we will not use this fact. We will only prove this invariance in the almost-homogeneous case; see Lemma 8.2.
@@
In this section we finish the proof of Theorem 1.4. We take $\Omega\subset\mathbf{Sb}(G)$ an almost-homogeneous domain, proper in $\mathbb{A}$.
Let $x\in\Omega$ and let $p_{0}\in\partial\Omega\cap\mathbf{J}^{-}(x)$ and $q_{0}\in\partial\Omega\cap\mathbf{J}^{+}(x)$ be two strongly $\mathscr{R}$-extremal points of $\partial\Omega$ given by Lemma 7.7. Then in particular $p_{0},q_{0}\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$ (Remark 7.8.(2)), so by Theorem 7.2
@@
$\mathsf{Aut}(\Omega)\leq\mathsf{Aut}(\mathbf{D}(p_{0},q_{0})).$
Since $\mathbf{D}(p_{0},q_{0})$ is proper, Lemma 4.4 implies that $\Omega=\mathbf{D}(p_{0},q_{0})$. This concludes the proof of Theorem 1.4.
@@ 9. Application: closed manifolds with proper development
    Let $\pi_{1}(M)$ be the fundamental group of $M$ and let $\mathsf{hol}:\pi_{1}(M)\rightarrow G$ be the holonomy of $M$, so that the map $\mathsf{dev}$ is $\mathsf{hol}$-equivariant. Let $\Omega=\mathsf{dev}(\widetilde{M})$. Since $M$ is closed, there is a compact fundamental domain $\mathscr{K}\subset\widetilde{M}$ intersecting every $\pi_{1}(M)$-orbit. By $\mathsf{hol}$-equivariance of $\mathsf{dev}$, the compact set $\mathsf{dev}(\mathscr{K})\subset\Omega$ intersects any $\mathsf{hol}(\pi_{1}(M))$-orbit. Hence $\Omega$ is almost-homogeneous and thus almost-homogeneous. Theorem 1.4 implies that $\Omega$ is a diamond and is $\left(H_{1},H_{2}\right)$-equivariantly diffeomorphic to the symmetric space ×$\mathbb{R}\times X_{L_{s}}$, where $H_{1},H_{2}$ are finite-index subgroupds of $\mathsf{Aut}(\Omega),\mathrm{Isom}\left(\mathbb{R}\times X_{L_{s}}\right)$ respectively. Up to taking a finite cover of $M$, we may assume that $H_{1}=\mathsf{Aut}(\Omega)$ and $H_{2}=\mathrm{Isom}(\mathbb{R}\times X_{L_{s}})$.